OpenStudy (anonymous):

Find the quotient of the complex numbers. Express your answer in trigonometric form. z1= 5(cos25Deg + i sin25Deg) z2= 2(cos80Deg + isin80Deg)

OpenStudy (anonymous):

Explain, please? (:

hartnn (hartnn):

when we divide complex numbers , we 1) divide the magnitude 2) subtract the angles.

hartnn (hartnn):

so, we need z1/z2 here, right ? what u get resultant magnitude after dividing the magnitudes of z1 and z2 ?

OpenStudy (anonymous):

so 5/2?

hartnn (hartnn):

yup, u got the magnitude as 5/2 now subtract the angles, 25-80 = ... ?

OpenStudy (anonymous):

-55 ? :o

hartnn (hartnn):

yes, to make it positive we can add or subtract 360 degrees, what do we do, add or subtract ?

OpenStudy (anonymous):

add.. :)

hartnn (hartnn):

correct, -55+360 = ... ?

OpenStudy (anonymous):

305 !

hartnn (hartnn):

you got it! :)

OpenStudy (anonymous):

Thanks you ! :D <3

OpenStudy (anonymous):

mind helping me with a few others? ^~^

hartnn (hartnn):

most welcome ^_^ i will be glad to :)

OpenStudy (anonymous):

OK, how do you simplify (2-3i)+ (5-6i) ? :)

hartnn (hartnn):

just addition of complex numbers., oh thats simple , add up "non-i" terms called real part, and "the i" part, called imaginary part like fo ex : (a+bi) - (c-di) = a-c + (b+d)i

OpenStudy (anonymous):

so 7+ -9i ?

hartnn (hartnn):

7-9i yup, absolutely correct! :D

OpenStudy (anonymous):

yay (: okay, and i is -1 right?

OpenStudy (anonymous):

or is correct just like that? .__.

hartnn (hartnn):

actually , its "square root" of -1 \(\large \color{green}{i=\sqrt{-1}}\)

OpenStudy (anonymous):

ohhh okay (: !

OpenStudy (anonymous):

got it !

OpenStudy (anonymous):

Find the product of the complex numbers. Express your answer in trigonometric form. z1= 7(cos 15Deg + i sin 15Deg) z2=2(cos 110Deg+ i sin 110Deg) ...? :o

hartnn (hartnn):

when we multiply complex numbers , we 1) multiply the magnitude 2) add the angles. give it a shot!

OpenStudy (anonymous):

so 14, 125 and 125 :D

OpenStudy (anonymous):

14(cos 125Deg +i sin 125 Deg)

hartnn (hartnn):

\(\huge \color{red}\checkmark\)

OpenStudy (anonymous):

:D <3

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[z1=5\left( \cos 25+\iota \sin 25 \right)=5 e ^{25\iota} \] \[z2=2\left( \cos 80+\iota \sin 80 \right)=2 e ^{80 \iota} \] \[\frac{ z1 }{z2 }=\frac{ 5e ^{25 \iota} }{2 e ^{80 \iota} }=\frac{ 5 }{2 }e ^{25 \iota-80 \iota}=\frac{ 5 }{2 } e ^{-55 \iota} =\frac{ 5 }{ 2 }e ^{\iota \left( -55+360 \right)}=\frac{ 5 }{2 }e ^{305 \iota} =\frac{ 5 }{2 } \left( \cos 305+\iota \sin 305 \right) \] \]

OpenStudy (anonymous):

Find the product of the complex numbers. Express your answer in trigonometric form.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Do we multiply the magnitude here too ?.. :o

hartnn (hartnn):

yup, try ?

OpenStudy (anonymous):

so.... 12

hartnn (hartnn):

yes and \(\pi/3+2\pi/5 =...?\)

OpenStudy (anonymous):

\[11\pi/15\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

D: idk

hartnn (hartnn):

yup, correct :)

OpenStudy (anonymous):

yay! (:

OpenStudy (anonymous):

When it asks what is the conjugate of 5+2i what does that mean?

hartnn (hartnn):

conjugate of \(\large a+bi\) is simply \(\large a-bi\) like conjugate of 3-4i will be 3+4i :)

OpenStudy (anonymous):

so.. 5-2i ?? :D

OpenStudy (anonymous):

the opposite (: !

hartnn (hartnn):

yes! you learn fast :D

OpenStudy (anonymous):

yay ! :D yeah, I just need a little push sometimes from nice and smart strangers like you x__x

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Find the quotient of z1 by z2. Express your answer in trigonometric form.

OpenStudy (anonymous):

So.... you multiply !! :D RIght? :)

hartnn (hartnn):

quotient means dividing... and what do we do to divide complex numbers ?

OpenStudy (anonymous):

3/7 ...

OpenStudy (anonymous):

:o and i'm not sure. .. .__. :o Lol

hartnn (hartnn):

my very first comment! when we divide complex numbers , we 1) divide the magnitude 2) subtract the angles. so 3/7 is correct what about the angle ?

OpenStudy (anonymous):

RIIIGHHHT. so \[-\pi/8\] ?

hartnn (hartnn):

yup! is that in any of the choices ?

OpenStudy (anonymous):

mhm (: